### v1

@@ 1   Introduction
Understanding and predicting traffic flow dynamics require accurate modeling of individual vehicle behavior, which serves as a foundation for both theoretical explorations and practical engineering applications in transportation systems. Among the various modeling approaches, car following (CF) models play a pivotal role by describing how drivers adjust their vehicle speed based on interactions with surrounding vehicles. These models provide essential tools for analyzing traffic stability, simulating diverse traffic scenarios, and optimizing intelligent transportation systems (ITS). Over the decades, several classical CF models have been proposed, including the GM model (Chandler et al., 1958), the GHR model (Gazis et al., 1961), Gipps’ model (Gipps, 1981), the Wiedemann model (Wiedemann and Reiter, 1992), the Optimal Velocity Model (OVM, (Bando et al., 1995)), the Full Velocity Difference Model (FVDM, (Jiang et al., 2001)), and Newell’s model (Newell, 2002). Each of these models has contributed to advancing our understanding of driver behavior and traffic flow, offering varying levels of realism, computational efficiency, and theoretical rigor.
@@
Figure 1: Number of papers published between 2000 and 2024 that use IDM for at least simulation or theoretical analysis (rather than just citing IDM). Blue and red bars represent relevant papers published in the target list of journals (see Appendix) and in all journals in the Web of Science, respectively.
@@
Among these, the Intelligent Driver Model (IDM), introduced by Treiber et al. (2000), has emerged as one of the most widely adopted CF models in traffic flow research and applications (see Fig. 1). IDM is recognized for its simplicity, computational efficiency, and strong physical interpretability. Unlike earlier models, IDM conceptualizes vehicles as intelligent agents, dynamically adjusting their acceleration based on nonlinear interactions with the leading vehicle, while incorporating factors such as desired speed, safe distance, and relative velocity. This design allows IDM to capture a wide range of traffic states, including free flow, congestion, and stop-and-go oscillations, and to model complex phenomena like phase transitions, hysteresis, and multi-stability. Its versatility has made IDM a benchmark in both theoretical traffic flow studies and practical ITS applications.
@@
Since its introduction in 2000, IDM has been extensively utilized in academic research and real-world applications. Table 1 indicates that the IDM has garnered 3458 citations, averaging 138.32 citations per year (CPY). It ranks first in both total citations and CPY among classic CF models. Moreover, compared to more general models of traffic flow, although IDM has not acquired the highest citations, its CPY always ranks first. It has facilitated advancements in understanding CF behavior, deriving fundamental diagrams, and analyzing traffic flow stability. Moreover, IDM has been integrated into widely used platforms such as SUMO, enabling its application in evaluating vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) communication technologies, as well as in testing advanced driver-assistance systems and autonomous vehicle control algorithms. However, despite its widespread adoption, IDM’s deterministic framework and simplified assumptions face challenges in capturing real-world traffic complexities, such as stochastic variability, driver heterogeneity, and the interactions between human-driven vehicles (HVs) and connected and autonomous vehicles (CAVs).
@@
Table 1: Citations of famous classic models included in the Web of Science Core as of May 2025. Wiedemann and Reiter (1992) is not included in the Web of Science, so the citation is collected in google scholar database.
{tblr}
width = colspec = Q[113]Q[48]Q[225]Q[319]Q[140]Q[90], cells = c, cell21 = r=10, cell121 = r=4, hline1,16 = -0.08em, hline2,12 = -, Classic Models & Citations Classic Models Title Publication Source Citations Per Year(CPY)
@@
(Treiber et al., 2000) Congested Traffic States in Empirical Observations and Microscopic Simulations Physical Review E 138.32
@@
(Bando et al., 1995) Dynamical model of traffic congestion and numerical simulation Physical Review E 87.60
@@
(Gipps, 1981) A behavioural CF model for computer simulation Transportation Research Part B 40.30
@@
(Bando et al., 1995) Full velocity difference model for a car following theory Physical Review E 54.92
@@
(Gazis et al., 1961) Nonlinear follow-the-leader models of traffic flow Operations Research 15.39
@@
(Chandler et al., 1958) Traffic dynamics: studies in car following Operations Research 13.93
881 Pipes Model(Pipes, 1953) An operational analysis of traffic dynamics Journal of applied physics 12.24
827 Newell Nonlinear Model(Newell, 1961) Nonlinear effects in the dynamics of car following Operations Research 12.92
@@
(Newell, 2002) A simplified car following theory: a lower order model Transportation Research Part B 27.87
@@
(Wiedemann and Reiter, 1992) Microscopic traffic simulation: the simulation system MISSION Project ICARUS (V1052) Final Report 11.45
Models in Transportation Science 3423 NaSch
(Nagel and Schreckenberg, 1992) A cellular automaton model for freeway traffic Journal De Physique 103.76
@@
(Lighthill and Whitham, 1955) On kinematic waves II. a theory of traffic flow on long crowded roads Proceedings of the Royal Society of London 47.41
@@
(Daganzo, 1994) The cell transmission model-a dynamic representation of highway traffic consistent with the hydrodynamic theory Transportation Research Part B 73.42
@@
(Vickrey, 1969) Congestion theory and transport investment American Economic Review 28.41
@@
{tblr}
hline1,15 = -0.08em, hline2 = -, rowodd = gray9 Abbreviations & Full names
@@ 2.1.1 Model Structure
where $a_{n}(t)$ represents the vehicle’s acceleration, $a$ is the maximum acceleration, $v_{n}(t)$ is the current velocity, $v_{0}$ is the desired velocity, $s_{n}(t)$ is the actual spacing, and $s^{*}_{n}(t)$ is the desired dynamic spacing. The exponent $\delta$ determines how quickly acceleration decreases as a vehicle approaches its desired speed. The term $\left(\frac{s^{*}}{s}\right)^{2}$ is used to describe the nonlinear sensitivity of drivers to changes in vehicle spacing, particularly as the actual spacing $s$ approaches the desired spacing $s^{*}$. The quadratic form of this term is closely aligned with Stevens’ power law, which states that the perceived intensity of a stimulus follows a power-law relationship with the actual stimulus. Specifically, the exponent $n=2$ in IDM reflects a balance between capturing drivers’ heightened sensitivity to small spacings and maintaining numerical stability in the model. While a linear form ($n=1$) would underestimate drivers’ rapid responses in critical situations, and a cubic form ($n=3$) would lead to excessive sensitivity, the quadratic term provides a reasonable approximation of how drivers respond nonlinearly to perceived distance changes, as suggested by Stevens’ law for psychophysical scaling (Park et al., 2001).
@@
where $s_{0}$ is the minimum spacing, $T$ is the desired time headway, $\Delta v_{n}(t)=v_{n}(t)-v_{n-1}(t)$ is the velocity difference to the preceding vehicle $n-1$, and $b$ is the comfortable deceleration. The model captures critical driving behaviors by decreasing acceleration as the vehicle approaches its desired speed and ensuring safety through dynamic spacing.
@@
The collision-free nature of the IDM has been analytically validated and confirmed through simulation studies (e.g., Treiber and Kesting (2013b)). This robust safety property makes IDM particularly well-suited for applications such as Advanced Driver Assistance Systems, where proactive collision avoidance is critical. On the other hand, it makes IDM (or any safe CF model) not suitable for evaluating driving safety.
@@
Moreover, as outlined in Equations (1) and (2), the IDM is a nonlinear CF model characterized by six parameters: $a$, $b$, $v_{0}$, $\delta$, $s_{0}$, and $T$. According to Treiber and Kesting (2013a), these parameters are intuitive, each is related to a specific driving regime: the desired speed $v_{0}$ is relevant for cruising in free-flow scenarios, the desired time gap $T$ pertains to steady-state CF, the minimum gap $s_{0}$ is associated with creeping and standing traffic, and the maximum acceleration $a$ and desired deceleration $b$ relate to non-steady traffic flow. The exponent $\delta$ governs this reduction: a larger $\delta$ delays the reduction in acceleration as the vehicle approaches the desired velocity. In the limit $\delta\to\infty$, the acceleration curve aligns with that of the Gipps model, while $\delta=1$ replicates the overly smooth acceleration behavior characteristic of the optimal velocity model. This design not only gives IDM strong physical meaning but also provides analytical value.
@@ 2.1.2 Model Philosophies
Hence, IDM is widely used not only because it has a clear physical meaning, but also because it was developed with some basic assumptions built on the modeling philosophy (Treiber and Kesting, 2013b):
@@ 2.1.3 Connections with Classic Models
The modeling philosophy of IDM is not an isolated concept but builds upon a rich history of CF theories. Many of its principles, including safety constraints, equilibrium distances, and smooth transitions between driving states, are inspired by or extend ideas from earlier CF models. To fully understand its motivations and distinctive characteristics, it is essential to examine its relationship to foundational models such as the GM model (Chandler et al., 1958), the OVM (Bando et al., 1995) and the Gipps model (Gipps, 1981). IDM synthesizes and extends their strengths while addressing their inherent limitations, resulting in a model that is both comprehensive and robust.
@@ 2.2.1 Numerical Schemes of the IDM
For numerical solution schemes of the IDM, Treiber and Kanagaraj (2015) compared four explicit integration methods: Euler method, ballistic update, Heun method (trapezoidal rule), and standard fourth-order Runge-Kutta (RK4). The study found that the ballistic scheme is the best for all practical purposes. Particularly, it is better than Runge-Kutta 4 (RK4) and Euler. The higher consistency order 4 of RK4 is of no use when lane changes or other discrete events are included reducing the overall consistency order to 1. While Euler has the same consistency order 1 as the ballistic scheme, the prefactor of the numerical error is much smaller. Moreover, the ballistic scheme with larger update time gaps could be considered as a special case of the human factor "limited attention span/action points" with an action (i.e., a changed acceleration) only occurring at each timestep. Most researchers and simulators, including SUMO, use this scheme.
From the parameter calibration perspective, Treiber and Kesting (2013a) explored methodologies for calibrating CF models based on IDM. They emphasized criteria for assessing model quality in addition to RMSE, including completeness, robustness, parameter orthogonality, and intuitive parameters aligned with plausible values. The study highlighted that global calibration based on gaps is more reliable than local calibration or those based on speeds or accelerations. Moreover, they found that a sampling rate of 1 Hz suffices, with data completeness and a minimum interval of 300 seconds being crucial for accurate calibration, while data smoothing had minimal impact. Sharma et al. (2019) found that there is no one-to-one correspondence between IDM parameters and driving regimes. Specifically, the acceleration behavior of the IDM driver in a specific state is controlled by the interaction of multiple IDM parameters.This highlights the importance of considering parameter interactions and dataset characteristics when interpreting IDM calibration results. Punzo et al. (2015) conducted a comprehensive analysis of IDM parameters, identifying the minimum time headway $T$ as the most influential parameter in determining model performance and stability. Sensitivity analysis revealed that $T$ explains the majority of performance variability, followed by the acceleration exponent $\alpha$ and maximum acceleration $a$. To balance model simplicity and performance, they proposed a simplified IDM configuration by fixing less influential parameters, achieving faster convergence while maintaining comparable accuracy. This aligns with other findings emphasizing the importance of prioritizing key parameters during calibration for robust model performance.
@@ 2.2.2 Traffic Stability
The interplay of these parameters determines the overall stability of traffic flow modeled by IDM. For more insights and mathematical formulations, readers are referred to Treiber and Kesting (2013b).
@@
Despite its practical utility, the IDM does exhibit limitations when compared to empirical observations. For instance, while IDM can reproduce certain instability phenomena, such as upstream-propagating convective instabilities and transitions between convective and absolute instabilities, its predictive accuracy for traffic fluctuations is limited due to inherent oversimplifications. In Jiang et al. (2014), experimental evidence shows that traffic oscillations exhibit a concave growth pattern, where the standard deviation of vehicle acceleration increases concavely as the disturbance propagates upstream in a platoon (see Fig. 2). This finding has been further validated using various datasets, including NGSIM data and controlled experiments, such as the 25- and 51-vehicle platoon experiments in Hefei (Jiang et al. (2015); Tian et al. (2016a); Tian et al. (2016c)). Together, these studies confirm that the concave growth of traffic oscillations is a universal phenomenon observed under different traffic conditions and experimental settings. In contrast, traditional CF models, such as the IDM, inherently predict a convex growth pattern of oscillations due to the linear instability of their steady-state solutions (Wang et al., 2020b). This convex growth highlights a critical limitation of IDM, as it fails to capture the empirically observed concave growth, which is essential for accurately modeling traffic flow dynamics.
Moreover, the IDM assumes a deterministic framework that overlooks the stochastic and heterogeneous nature of driver behavior, which plays a significant role in traffic stability. Further insights from Tian et al. (2019) suggest that the fundamental mechanism driving traffic fluctuations lies in the competition between speed adaptation and stochastic effects, rather than being solely attributed to reaction delay as assumed in the IDM. Their analysis of CF behavior in a 25-car platoon experiment shows that reaction delay has a negligible effect on fluctuation growth, while the interplay of speed adaptation and stochastic disturbance dominates the dynamics. This evidence challenges IDM’s core assumptions and indicates that its deterministic nature limits its ability to accurately model real-world instability.
Figure 2: Simulation results (solid lines) of the standard deviation $\sigma_{v}$ of the time series of the velocity of each car of IDM from Jiang et al. (2014). The curves represent the simulation results while the markers such as circles represent the experimental results. The different colors represent different experiment runs in which the velocity of leading vehicle $v_{l}$ is 7,15,30,40 and 50 km/h.

### v2

@@ 1  Introduction
Understanding and predicting traffic flow dynamics require accurate modeling of individual vehicle behavior, which serves as a foundation for both theoretical explorations and practical engineering applications in transportation systems. Among the various modeling approaches, car following (CF) models play a pivotal role by describing how drivers adjust their vehicle speed based on interactions with surrounding vehicles. These models provide essential tools for analyzing traffic stability, simulating diverse traffic scenarios, and optimizing intelligent transportation systems (ITS). Over the decades, several classical CF models have been proposed, including the GM model (chandlerTrafficDynamicsStudiesCar1958), the GHR model (gazisNonlinearFollowLeaderModelsTraffic1961), Gipps’ model (gippsBehaviouralCarfollowingModelComputer1981), the Wiedemann model (wiedemann1992microscopic), the Optimal Velocity Model (OVM, (bandoDynamicalModelTrafficCongestion1995)), the Full Velocity Difference Model (FVDM, (jiangFullVelocityDifferenceModel2001)), and Newell’s model (newellSimplifiedCarfollowingTheoryLower2002). Each of these models has contributed to advancing our understanding of driver behavior and traffic flow, offering varying levels of realism, computational efficiency, and theoretical rigor.
@@
Figure 1: The red bars indicate the number of papers citing IDM in the Web of Science between 2000 and 2024. The blue bars indicate relevant papers published in the target list of journals (see Appendix) that at least use IDM for simulation or theoretical analysis rather than just citing it.
@@
Among these, the Intelligent Driver Model (IDM), introduced by treiberCongestedTrafficStatesEmpirical2000, has emerged as one of the most widely adopted CF models in traffic flow research and applications (see Fig. 1). IDM is recognized for its simplicity, computational efficiency, and strong physical interpretability. Unlike earlier models, IDM conceptualizes vehicles as intelligent agents, dynamically adjusting their acceleration based on nonlinear interactions with the leading vehicle, while incorporating factors such as desired speed, safe distance, and relative velocity. This design allows IDM to capture a wide range of traffic states, including free flow, congestion, and stop-and-go oscillations, and to model complex phenomena like phase transitions, hysteresis, and multi-stability. Its versatility has made IDM a benchmark in both theoretical traffic flow studies and practical ITS applications.
@@
Since its introduction in 2000, IDM has been extensively utilized in academic research and real-world applications, see Fig. 1. Table 1 indicates that the IDM has garnered 3458 citations, averaging 138.32 citations per year (CPY). It ranks first in both total citations and CPY among not only classic CF models but also more general models of transportation science. It has facilitated advancements in understanding CF behavior, deriving fundamental diagrams, and analyzing traffic flow stability. Moreover, IDM has been integrated into widely used platforms such as SUMO, enabling its application in evaluating vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) communication technologies, as well as in testing advanced driver-assistance systems and autonomous vehicle control algorithms. However, despite its widespread adoption, IDM’s deterministic framework and simplified assumptions face challenges in capturing real-world traffic complexities, such as stochastic variability, driver heterogeneity, and the interactions between human-driven vehicles (HVs) and connected and autonomous vehicles (CAVs).
@@
Table 1: Citations of famous classic models included in the Web of Science Core as of May 2025. wiedemann1992microscopic is not included in the Web of Science, so the citation is collected in google scholar database.
Classic Models Citations Classic Models Title Publication Source Citations Per Year(CPY)
@@
(treiberCongestedTrafficStatesEmpirical2000) Congested Traffic States in Empirical Observations and Microscopic Simulations Physical Review E 138.32
@@
(bandoDynamicalModelTrafficCongestion1995) Dynamical model of traffic congestion and numerical simulation Physical Review E 87.60
@@
(gippsBehaviouralCarfollowingModelComputer1981) A behavioural CF model for computer simulation Transportation Research Part B 40.30
@@
(bandoDynamicalModelTrafficCongestion1995) Full velocity difference model for a car following theory Physical Review E 54.92
@@
(gazisNonlinearFollowLeaderModelsTraffic1961) Nonlinear follow-the-leader models of traffic flow Operations Research 15.39
@@
(chandlerTrafficDynamicsStudiesCar1958) Traffic dynamics: studies in car following Operations Research 13.93
881 Pipes Model(pipesOperationalAnalysisTrafficDynamics1953) An operational analysis of traffic dynamics Journal of applied physics 12.24
827 Newell Nonlinear Model(newellNonlinearEffectsDynamicsCar1961a) Nonlinear effects in the dynamics of car following Operations Research 12.92
@@
(newellSimplifiedCarfollowingTheoryLower2002) A simplified car following theory: a lower order model Transportation Research Part B 27.87
@@
(wiedemann1992microscopic) Microscopic traffic simulation: the simulation system MISSION Project ICARUS (V1052) Final Report 11.45
More General Models of Transportation Science 3423 NaSch
(nagelCellularAutomatonModelFreeway1992a) A cellular automaton model for freeway traffic Journal De Physique 103.76
@@
(lighthillKinematicWavesIITheory1955) On kinematic waves II. a theory of traffic flow on long crowded roads Proceedings of the Royal Society of London 47.41
@@
(daganzoCellTransmissionModelDynamic1994) The cell transmission model-a dynamic representation of highway traffic consistent with the hydrodynamic theory Transportation Research Part B 73.42
@@
(vickreyCongestionTheoryTransportInvestment1969) Congestion theory and transport investment American Economic Review 28.41
@@
Abbreviations Full names
@@ 2.1.1 Model Structure
where $a_{n}(t)$ represents the vehicle’s acceleration, $a$ is the maximum acceleration, $v_{n}(t)$ is the current velocity, $v_{0}$ is the desired velocity, $s_{n}(t)$ is the actual spacing, and $s^{*}_{n}(t)$ is the desired dynamic spacing. The exponent $\delta$ determines how quickly acceleration decreases as a vehicle approaches its desired speed. The term $\left(\frac{s^{*}}{s}\right)^{2}$ is used to describe the nonlinear sensitivity of drivers to changes in vehicle spacing, particularly as the actual spacing $s$ approaches the desired spacing $s^{*}$. The quadratic form of this term is closely aligned with Stevens’ power law, which states that the perceived intensity of a stimulus follows a power-law relationship with the actual stimulus. Specifically, the exponent $n=2$ in IDM reflects a balance between capturing drivers’ heightened sensitivity to small spacings and maintaining numerical stability in the model. While a linear form ($n=1$) would underestimate drivers’ rapid responses in critical situations, and a cubic form ($n=3$) would lead to excessive sensitivity, the quadratic term provides a reasonable approximation of how drivers respond nonlinearly to perceived distance changes, as suggested by Stevens’ law for psychophysical scaling (parkDriversCharacteristicsPerceptionLead2001).
@@
where $s_{0}$ is the minimum spacing, $T$ is the desired time headway, $\Delta v_{n}(t)=v_{n}(t)-v_{n-1}(t)$ is the velocity difference to the preceding vehicle n1n-1, and $b$ is the comfortable deceleration. The model captures critical driving behaviors by decreasing acceleration as the vehicle approaches its desired speed and ensuring safety through dynamic spacing.
@@
The collision-free nature of the IDM has been analytically validated and confirmed through simulation studies (e.g., treiberTrafficFlowDynamics2013). This robust safety property makes IDM particularly well-suited for applications such as Advanced Driver Assistance Systems, where proactive collision avoidance is critical. On the other hand, it makes IDM (or any safe CF model) not suitable for evaluating driving safety.
@@
Moreover, as outlined in Equations (1) and (2), the IDM is a nonlinear CF model characterized by six parameters: $a$, $b$, $v_{0}$, $\delta$, $s_{0}$, and $T$. According to treiberMicroscopicCalibrationValidationCarFollowing2013, these parameters are intuitive, each is related to a specific driving regime: the desired speed $v_{0}$ is relevant for cruising in free-flow scenarios, the desired time gap $T$ pertains to steady-state CF, the minimum gap $s_{0}$ is associated with creeping and standing traffic, and the maximum acceleration $a$ and desired deceleration $b$ relate to non-steady traffic flow. The exponent $\delta$ governs this reduction: a larger $\delta$ delays the reduction in acceleration as the vehicle approaches the desired velocity. In the limit $\delta\to\infty$, the acceleration curve aligns with that of the Gipps model, while $\delta=1$ replicates the overly smooth acceleration behavior characteristic of the optimal velocity model. This design not only gives IDM strong physical meaning but also provides analytical value.
@@ 2.1.2 Model Philosophies
Hence, IDM is widely used not only because it has a clear physical meaning, but also because it was developed with some basic assumptions built on the modeling philosophy (treiberTrafficFlowDynamics2013):
@@ 2.1.3 Connections with Classic Models
The modeling philosophy of IDM is not an isolated concept but builds upon a rich history of CF theories. Many of its principles, including safety constraints, equilibrium distances, and smooth transitions between driving states, are inspired by or extend ideas from earlier CF models. To fully understand its motivations and distinctive characteristics, it is essential to examine its relationship to foundational models such as the GM model (chandlerTrafficDynamicsStudiesCar1958), the OVM (bandoDynamicalModelTrafficCongestion1995) and the Gipps model (gippsBehaviouralCarfollowingModelComputer1981). IDM synthesizes and extends their strengths while addressing their inherent limitations, resulting in a model that is both comprehensive and robust.
@@ 2.2.1 Numerical Schemes of the IDM
For numerical solution schemes of the IDM, treiberComparingNumericalIntegrationSchemes2015 compared four explicit integration methods: Euler method, ballistic update, Heun method (trapezoidal rule), and standard fourth-order Runge-Kutta (RK4). The study found that the ballistic scheme is the best for all practical purposes. Particularly, it is better than Runge-Kutta 4 (RK4) and Euler. The higher consistency order 4 of RK4 is of no use when lane changes or other discrete events are included reducing the overall consistency order to 1. While Euler has the same consistency order 1 as the ballistic scheme, the prefactor of the numerical error is much smaller. Moreover, the ballistic scheme with larger update time gaps could be considered as a special case of the human factor "limited attention span/action points" with an action (i.e., a changed acceleration) only occurring at each timestep. Most researchers and simulators, including SUMO, use this scheme.
From the parameter calibration perspective, treiberMicroscopicCalibrationValidationCarFollowing2013 explored methodologies for calibrating CF models based on IDM. They emphasized criteria for assessing model quality in addition to RMSE, including completeness, robustness, parameter orthogonality, and intuitive parameters aligned with plausible values. The study highlighted that global calibration based on gaps is more reliable than local calibration or those based on speeds or accelerations. Moreover, they found that a sampling rate of 1 Hz suffices, with data completeness and a minimum interval of 300 seconds being crucial for accurate calibration, while data smoothing had minimal impact. sharmaMoreAlwaysBetterImpact2019b found that there is no one-to-one correspondence between IDM parameters and driving regimes. Specifically, the acceleration behavior of the IDM driver in a specific state is controlled by the interaction of multiple IDM parameters.This highlights the importance of considering parameter interactions and dataset characteristics when interpreting IDM calibration results. punzo_we_2015 conducted a comprehensive analysis of IDM parameters, identifying the minimum time headway $T$ as the most influential parameter in determining model performance and stability. Sensitivity analysis revealed that $T$ explains the majority of performance variability, followed by the acceleration exponent $\alpha$ and maximum acceleration $a$. To balance model simplicity and performance, they proposed a simplified IDM configuration by fixing less influential parameters, achieving faster convergence while maintaining comparable accuracy. This aligns with other findings emphasizing the importance of prioritizing key parameters during calibration for robust model performance.
@@ 2.2.2 Traffic Stability
The interplay of these parameters determines the overall stability of traffic flow modeled by IDM. For more insights and mathematical formulations, readers are referred to treiberTrafficFlowDynamics2013.
@@
Despite its practical utility, the IDM does exhibit limitations when compared to empirical observations. For instance, while IDM can reproduce certain instability phenomena, such as upstream-propagating convective instabilities and transitions between convective and absolute instabilities, its predictive accuracy for traffic fluctuations is limited due to inherent oversimplifications. In jiangTrafficExperimentRevealsNature2014, experimental evidence shows that traffic oscillations exhibit a concave growth pattern, where the standard deviation of vehicle acceleration increases concavely as the disturbance propagates upstream in a platoon (see Fig. 2). This finding has been further validated using various datasets, including NGSIM data and controlled experiments, such as the 25- and 51-vehicle platoon experiments in Hefei (jiangExperimentalFeaturesCarfollowingBehavior2015; tianEmpiricalAnalysisSimulationConcave2016; tianCellularAutomatonModelSimulating2016). Together, these studies confirm that the concave growth of traffic oscillations is a universal phenomenon observed under different traffic conditions and experimental settings. In contrast, traditional CF models, such as the IDM, inherently predict a convex growth pattern of oscillations due to the linear instability of their steady-state solutions (wangStabilityAnalysisStochasticLinear2020). This convex growth highlights a critical limitation of IDM, as it fails to capture the empirically observed concave growth, which is essential for accurately modeling traffic flow dynamics.
Moreover, the IDM assumes a deterministic framework that overlooks the stochastic and heterogeneous nature of driver behavior, which plays a significant role in traffic stability. Further insights from tianRoleSpeedAdaptationSpacing2019 suggest that the fundamental mechanism driving traffic fluctuations lies in the competition between speed adaptation and stochastic effects, rather than being solely attributed to reaction delay as assumed in the IDM. Their analysis of CF behavior in a 25-car platoon experiment shows that reaction delay has a negligible effect on fluctuation growth, while the interplay of speed adaptation and stochastic disturbance dominates the dynamics. This evidence challenges IDM’s core assumptions and indicates that its deterministic nature limits its ability to accurately model real-world instability.
Figure 2: Simulation results (solid lines) of the standard deviation $\sigma_{v}$ of the time series of the velocity of each car of IDM from jiangTrafficExperimentRevealsNature2014. The curves represent the simulation results while the markers such as circles represent the experimental results. The different colors represent different experiment runs in which the velocity of leading vehicle $v_{l}$ is 7,15,30,40 and 50 km/h.